OpenStudy (anonymous):

The value of x satisfying \[\Large \frac{ 6x+2a+3b+c }{ 6x+2a-3b-c }=\frac{ 2x+6a+b+3c }{ 2x+6a-b-3c }\]

OpenStudy (anonymous):

apply componendo and dividendo rule both sides to simplify and then further simplify

OpenStudy (ikram002p):

\( \huge \frac{ 6x+2a+3b+c }{ 6x+2a-3b-c }=\frac{ 2x+6a+b+3c }{ 2x+6a-b-3c } \) there is two cases :- i) \(6x+2a+3b+c =2x+6a+b+3c=0 \) ii) or , there is k s.t \(2x+6a+b+3c = k(2x+6a+b+3c )\) \(6x+2a-3b-c =k(2x+6a-b-3c )\)

OpenStudy (ikram002p):

i made a typo , in second case/first line 6x+2a+3b+ c = k(2x+6a+b+3c )

OpenStudy (anonymous):

(6x+2a)/(3b+c) = (2x+6a)/(b+3c) 2*(3x+a) /(3b+c) = 2(x+3a)/(b+3c) (3x+a)/(x+3a) = (3b+c)/(b+3c) (3x+a)/(x+3a) - 3 = (3b+c)/(b+3c) -3 (3x+a-3x-9a)/(x+3a) = (3b+c-3b -9c) /(b+3c) -8a/(x+3a) = -8c/(b+3c) (x+3a)/a=(b+3c)/c x/a +3 =b/c +3 x/a = b/c x=ab/c

ganeshie8 (ganeshie8):

\[ \frac{ 6x+2a+3b+c }{ 6x+2a-3b-c }=\frac{ 2x+6a+b+3c }{ 2x+6a-b-3c } \stackrel{\color{red}{C\&D}}{\iff} \frac{ 3x+a}{ 3b+c }=\frac{ 3a+x }{ 3c+b } \]

ganeshie8 (ganeshie8):

+1 for C&D :)

OpenStudy (ikram002p):

what is C&D ??

ganeshie8 (ganeshie8):

Componendo and Dividendo : \[\large \dfrac{a}{b} = \dfrac{c}{d} \iff \dfrac{a+b}{a-b} = \dfrac{c+d}{c-d}\] http://en.wikipedia.org/wiki/Componendo_and_dividendo

ganeshie8 (ganeshie8):

it just follows from adding 1 to both sides, but really more powerful than it appears when you look at it first time !

OpenStudy (ikram002p):

hehe cool xD ik this one though if 0<c , a<b then a-c<b+c hehe

ganeshie8 (ganeshie8):

true, but hows that related here ?

OpenStudy (ikram002p):

hehe nothing xD i just remmembered it :P

ganeshie8 (ganeshie8):

that wiki link has a short+nice proof :)

OpenStudy (ikram002p):

i read it , was cute to know ;) nice addition to my day lol

OpenStudy (anonymous):

Sorry about late reply i got it nice

OpenStudy (kainui):

So simple, and yet I don't know this. Cool haha, this is just the kind of thing that shows up with triangles all the time and sometimes chemistry.